Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 17:09 Saturday 10 April 2021 (51382) Print this report
Operator Report

The shift was dedicated to the planned OptChar Activity, FP cavities characterization, carried out by Guo, Capocasa, and Degallaix remotely.
To allow an initial measurement, I manually unlocked the Mode Cleaner from 8:30 UTC to 8:33 UTC, the system went back online without further adjustment. Also, as preliminary action, the Flip Mirrors dedicated to ALS at the Terminal Buildings were moved Up for the whole shift.
The following actions were focused on the lock of the infrared cavities. With only one set of mirrors aligned at times, we relock the arms, unlock them manually and misalign the End Mirrors, waiting after every action a specific time. Also, from around 12:00 UTC to 12:30 UTC, we set the ITF with only the NI aligned and moved the mirror to maximize the power on B4 as a reference.
The Activity was concluded at 13:40 UTC, infrared cavities left locked.

Comments to this report:
guo - 13:24 Monday 12 April 2021 (51390) Print this report

The GPS time of all the activities can be found in the attached file.

Non-image files attached to this comment
Search Help